OpenStudy (anonymous):

A megabyte is two to the twentieth bytes. A gigabyte is two to the thirtieth bytes. Using the exponent rule for division, find how many megabytes are in a gigabyte.

OpenStudy (whpalmer4):

Exponent rule for division of numbers with identical bases is \[\frac{x^n}{x^m} = x^{n-m}\]

OpenStudy (anonymous):

\[megabyte = 2^{20} , gigabyte = 2^{30}\]

OpenStudy (whpalmer4):

\[\frac{2^{30}}{2^{20}} = \]